# Evaluating Sample Utility for Data Selection by Mimicking Model Weights

### v1

@@ 5.1 Simulating Mislabeled Samples
Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic consistently improves CLIP model performance.
((a))
((b))
Figure 4: Mimic scores offer precise dataset quality estimation: Left figure (a) computes the portion of retained samples to evaluate dataset quality in simulation experiments. Right figure (b) takes the average on mimic scores in each filtered dataset for estimation. [S/M] denotes the scale of DataComp dataset—small [S] or medium [M]—that the filter is designed on.
Stage 2 Results.
Next, we demonstrate Grad-Mimic’s capability to detect mislabeled samples. We use F1-scores to report the detection results in Table 2. Grad-Mimic accurately identifies mislabeled samples across all datasets and demonstrates robust performance when facing different noise levels.
We use CIFAR100 dataset as an example and visualize mimic score distribution at each training step in Figure 3. In this dataset, half of the labels are flipped. We see their distributions clearly separate into two groups, confirming that mimic scores can serve as informative signals to identify samples. Despite variations over training, these dynamics can be effectively captured in our aggregation step.
We use an ensemble filter from each dataset created by GMM clustering to estimate dataset quality using the proportion of retained data. The results, displayed in Figure 4, highly correspond to the presence of label noise, with a Pearson correlation of -0.903. This demonstrates mimic score’s effectiveness for dataset quality estimation.
5.2 Data Selection in Large-scale Web Datasets
Setups.
We evaluate Grad-Mimic in a more challenging setting using million-scale web-crawled datasets. We use the small- and medium-scale DataComp datasets [16], which contain approximately 10 million and 100 million image-caption pairs [^1] We identified that some URLs provided by DataComp dataset are now broken. This means that our results might not be comparable to previous approaches on the DataComp Leaderboard. See here for details., respectively, for CLIP model pretraining [2]. We follow the training setup from DataComp [53, 16] and use publicly available pretrained CLIP weights as our reference  [^2] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K. This pretrained model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us. It serves as a proxy to the ideal reference point for scenarios where training such powerful models is infeasible. We target to mimic the final MLP layer in the text and image encoders respectively. We evaluate model performance using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks. Additional training details can be found in Appendix C.2.
In Stage 2, we use top-$k$ percent selection method to create the final filter, subsequently training a CLIP model on the curated dataset to validate filter’s effectiveness.
@@ 5.2 Data Selection in Large-scale Web Datasets
We anticipate mimic scores help large-scale training focus on high-value samples, while the derived scores will be instrumental in automating effective data curation.
@@
In the first stage, we compare Grad-Mimic to vanilla training, where each sample contributes in the weight update equally. After training, we compare our mimic score-based filter against the following methods: (1) No Filtering: using the entire training pool, (2) Basic Filtering [16]: selecting samples based on predefined criteria like caption length, image size, and caption language—representing a human-designed filter, (3) CLIP Score [17]: selecting top-$k$ percent samples based on embedding alignment between images and captions. Scores are computed by OpenAI’s CLIP ViT-B/32 model pretrained on over 400 million samples [2]. We see CLIP score as a strong baseline, which is prebuilt through heavy training using hundreds of millions of samples.
@@
Table 3 presents pretraining results for CLIP models on both dataset scales. Grad-Mimic consistently outperforms standard training across all the temperature settings. Interestingly, we find that mimicking the last layer of the image encoder yields more performance gains compared to targeting the text encoder. These results highlight Grad-Mimic’s success in using pretrained model weights as a guide to prioritize high-value samples to learn.
@@
CLIP Score $\setminus$ Mimic Score Bottom 15% 2.7M 0.047 0.053 0.174 0.106 0.163
Small CLIP Score $\setminus$ Mimic Score Bottom 20% 2.6M 0.047 0.052 0.173 0.113 0.164
@@
CLIP Score $\setminus$ Mimic Score Bottom 5% 28.1M 0.270 0.227 0.339 0.246 0.323
CLIP Score $\setminus$ Mimic Score Bottom 10% 27.7M 0.274 0.226 0.334 0.243 0.323
Medium CLIP Score $\setminus$ Mimic Score Bottom 30% 25.3M 0.264 0.223 0.330 0.244 0.322
@@
Stage 2 Results.
We use the derived mimic scores from our best-performing model in Stage 1 to design filters and evaluate their effectiveness. Results are presented in Table 4. Grad-Mimic successfully automates the selection process, outperforming the basic filtering strategy at both dataset scales. Additionally, we remove less useful samples identified by Grad-Mimic to improve CLIP score-based filter (B/32 Top 30%). This complementary approach not only enhances model performance but also improves data efficiency by reducing the training set size, specifically cutting down 4.7 million samples in the medium-scale dataset.
Lastly, we gather mimic scores from datasets that are curated by various CLIP score-based filters. Besides, we include the top-ranked approach during our time of experimentation [54], which uses an orthogonal approach based on object detection models. By calculating the average mimic score in each curated dataset, we compare with their corresponding performance gains. The results, shown in Figure 4, reveal a strong alignment between performance gains and average mimic scores, with a Pearson correlation of 0.958, demonstrating that mimic scores can serve as a reliable metric for estimating dataset quality.
Figure 5: The filtering logic behind mimic score aligns with handcrafted heuristics, particularly in caption length. Low-value samples (in the left figure) tend to have captions that are either too short or exceed the maximum token length, i.e., 77.
@@
Figure 6: Mimic score distribution and its relationship with CLIP score.
@@
We analyze high- and low-value samples ranked by mimic score. In Figure 5, we compare the token usage distribution in captions. We find that bottom-ranked samples often carry captions that are either too short (one or two words) or excessively long (exceeding the maximum token limit in CLIP’s tokenizer), unlike top-ranked samples. Moreover, mimic score-based filter effectively identifies low-value samples, such as Captcha-like images in the small-scale dataset and empty or visually incoherent images in the medium-scale dataset (see examples shown in Figure 1). These filtering patterns align with human intuition but are automatically captured by Grad-Mimic framework.
We select the top 30% and bottom 30% of samples ranked by their mimic scores and visualize their distribution along with corresponding CLIP scores. In Figure 6, we observe high-value samples distributed towards higher CLIP scores, while low-value samples are more concentrated in lower CLIP score ranges (half of the samples are scored less than 0.2). Furthermore, we calculate their correlation using randomly selected samples. For a better visualization, we normalize both scores into the range 0 to 1. We see a positive correlation between them, supporting the reliability of mimic scores in identifying high-quality samples.
@@
We explore an interesting application of mimic score: predicting the pretraining dataset. Specifically, we ask: how accurately can Grad-Mimic check whether a given sample was used to train a reference model? We test this hypothesis on the small-scale DataComp dataset. We first apply various pre-built filters to curate datasets and train CLIP models on each. Then, we use their final layer weights as the reference and apply Grad-Mimic to train on the entire data pool. Our goal is to see whether the top-ranked samples (matched in size to the curated datasets) identified by mimic scores appear in their curated dataset.
We evaluate our approach against random selection using Jaccard similarity and percentage of overlap. As shown in Table 5, mimic score-based selection identifies more samples used to train the reference model compared to random sampling. We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset by simply mimicking the final layer weights without direct access to their filtering steps.
@@ 6 Conclusion
We introduce a new data quality metric called Mimic Score, designed to assess samples’ utility in the training process. We quantify samples by measuring the alignment of per-sample gradient and the vector toward a pretrained reference model. Building on Mimic Score, we develop Grad-Mimic to automate data selection process. Empirically, we demonstrate that Grad-Mimic effectively prioritizes identified samples to learn, resulting in enhanced model performance. Moreover, Mimic Score can serve as a reliable metric for creating effective filters and estimating dataset quality. Developed filters outperform human-designed ones and exhibit a strong correlation with performance gains.
@@ 7 Acknowledgments
We thank Kwei-Herng Lai, Prasad Gabbur, Meng Cao, Ryan Jia, Russ Webb, Vedant Joshi, Albert Ge, and John Cooper for their helpful feedback and valuable discussion.
Appendix A Discussion
A.1 Broader Impacts and Limitations.
The importance of data selection has grown in building foundation models. We demonstrate that Grad-Mimic can effectively identify and exclude noisy or unhelpful samples. This leads to reduced computational and resource costs. Besides, Grad-Mimic provides the second benefit by enabling models to focus their training on high-value samples, which contributes to performance improvements.
We do not foresee explicit negative impacts arising from Grad-Mimic. However, if the reference model used for sample selection is unreliable, there remains a risk of refining suboptimal datasets. To address this, incorporating human expert-designed filtering rules may be necessary to ensure the quality of curated datasets.
Appendix B Grad-Mimic Algorithm
We summarize the algorithm steps in Grad-Mimic for computing mimic scores and updating model weights. Our algorithm supports various training settings, such as supervised learning and self-supervised learning. We have validated both training scenarios through simulation experiments (Sec. 5.1) and by training on web-crawled datasets with CLIP models (Sec.5.2). Moreover, these experiments evaluate Grad-Mimic performance across three different training configurations: full-parameter fine-tuning, linear probing, and training models from scratch.
@@ Appendix B Grad-Mimic Algorithm
samples’ mimic scores $\overline{m}_{i,t}$, trained model $\theta_{T}$
@@
     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright compute the vector as guide
@@
         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright compute per-sample gradients
         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|_{2}}$triangleright compute mimic scores
@@
     $\overline{m}_{:,t}=\frac{e^{-m_{:,t}/\tau}}{\sum_{j=1}^{b}e^{-m_{j,t}/\tau}}$triangleright normalize mimic scores
@@
Appendix C Experimental Details
We provide more details about our training setups, and computational resources.
C.1 Simulation Experiments
In the experiments of simulating mislabeled samples (Sec. 5.1), we fine-tune ViT-B/16 models pretrained on the ImageNet-21k dataset [55]. Each dataset is trained with a batch size of 32, a learning rate of 1e-4, the AdamW optimizer, and for 5 epochs. We evaluate Grad-Mimic using various temperature values $\tau$ (1.0, 0.9, 0.8, 0.7, 0.6, and 0.5) and report the performance at $\tau=0.5$ in the main paper. Results for other temperature values are presented in Table 6. To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets using identical configurations but with different random seeds for weight initialization. These experiments were performed on a Nvidia Tesla A100.
@@ C.1 Simulation Experiments
Table 6: Stage 1 Results in Simulation Experiment with Different Temperature: Lower temperature makes the model focus more on training with high-value samples, resulting in a higher testing accuracy.
C.2 DataComp Experiments
We adopt the training setup from DataComp [53, 16]. CLIP models are trained from scratch using a contrastive objective over image-caption pairs. Each model is trained for 5 epochs with a batch size of 4096. The total number of seen samples is 12.8M for the small-scale dataset and 128M for the medium-scale dataset. These experiments ran on 8 Nvidia Tesla A100s.
We test Grad-Mimic using different parts of weights in the reference model, specifically the final MLP layer in the image and text encoders. The layers used are “visual.transformer.resblocks.11.mlp.c_fc.weight” and “text.transformer.resblocks.11.mlp.c_fc.weight”. Grad-Mimic’s performance is evaluated using 38 diverse downstream tasks [16] with various temperature values $\tau$ (0.03, 0.05, 0.07, 0.3, and 0.5). Results are presented in Table 3.
Appendix D Ablation Studies
@@
We present Grad-Mimic results with different temperature values and compare them to baseline methods (SGD and Grad-Match [24]). The results in the simulation experiments are presented in Table 6. We set the noise level to 0.3 and fine-tune ViT model under linear probing configuration. Grad-Mimic outperforms the baseline methods across all temperature settings. Lower temperatures yield better testing accuracy, as they normalize mimic scores in the way that encourages the model to focus more on high-value samples during training.

### v2

@@ 5.2 Data Selection in Large-scale Web Datasets
Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic yields higher CLIP model performance. Full table with various temperature settings is placed in Table 6 in Appendix  C.
@@
We anticipate mimic scores help large-scale training focus on high-value samples and can serve as instrumental metric for automating effective data curation.
@@
In the first stage, we compare Grad-Mimic to vanilla training, where each sample contributes in the weight update equally. After training, we compare our mimic score-based filter against the following methods: (1) No Filtering: using the entire training pool, (2) Basic Filtering [17]: selecting samples based on predefined criteria like caption length, image size, and caption language—representing a human-designed filter, (3) CLIP Score [18]: selecting top-$k$ percent samples based on embedding alignment between images and captions. Scores are computed by OpenAI’s CLIP ViT-B/32 model pretrained on over 400 million samples [2]. We see CLIP score as a strong baseline, as it is prebuilt through heavy training using hundreds of millions of samples.
@@
Table 3 presents pretraining results for CLIP models on both dataset scales. Grad-Mimic consistently outperforms standard training across all temperature settings. Full results using different temperatures are provided in Appendix C. Interestingly, we find that mimicking the last layer of the image encoder yields more performance gains compared to targeting the text encoder. These results highlight Grad-Mimic’s success in using online pretrained weights as a reference to prioritize high-value samples for learning.
@@
CLIP Score $\setminus$ Mimic Score Bottom 15% 2.7M (0.3M $\downarrow$) 0.047 0.053 0.174 0.106 0.163
Small CLIP Score $\setminus$ Mimic Score Bottom 20% 2.6M (0.4M $\downarrow$) 0.047 0.052 0.173 0.113 0.164
@@
CLIP Score $\setminus$ Mimic Score Bottom 5% 28.1M (1.9M $\downarrow$) 0.270 0.227 0.339 0.246 0.323
CLIP Score $\setminus$ Mimic Score Bottom 10% 27.7M (2.3M $\downarrow$) 0.274 0.226 0.334 0.243 0.323
Medium CLIP Score $\setminus$ Mimic Score Bottom 30% 25.3M (4.7M $\downarrow$) 0.264 0.223 0.330 0.244 0.322
@@
Figure 4: Mimic scores offer precise dataset quality estimation: Left figure computes the portion of retained samples to evaluate dataset quality in simulation experiments. Right figure takes the average on mimic scores in each filtered dataset for estimation. [S/M] denotes the scale of DataComp dataset—small [S] or medium [M]—that the filter is designed on.
Stage 2 Results.
We use the derived mimic scores from our best-performing model in Stage 1 to design filters and evaluate their effectiveness. Results are presented in Table 4. Grad-Mimic automates the selection, outperforming basic filtering strategy at both scales. Additionally, we remove low-value samples identified by mimic score to improve CLIP score-based filter (B/32 Top 30%). This complementary approach not only enhances model performance but also improves data efficiency by reducing the training set size, specifically cutting down 4.7 million samples in the medium-scale dataset.
Lastly, we gather mimic scores from datasets that are curated by various CLIP score-based filters. We include the top-ranked approach during our time of experimentation [50], which uses an orthogonal approach based on object detection models. We average samples’ mimic scores in each curated dataset and compare with their corresponding performance gains. The results, shown in Figure 4, reveal a positive alignment with a Pearson correlation of 0.958, demonstrating mimic score as a reliable metric for dataset quality estimation.
@@
We analyze high-value and low-value samples ranked by mimic score. In Figure 5, we compare token usage in captions. We find that bottom-ranked samples often carry captions that are either too short (1-2 words) or excessively long (over the maximum token limit), unlike top-ranked samples. Moreover, mimic score-based filter effectively identifies low-value samples, such as Captcha-like images in the small-scale dataset and empty or incoherent images in the medium-scale dataset (see examples in Figure 1). These filtering patterns align with human intuition but are automatically captured by Grad-Mimic.
We visualize the top and bottom 30% of samples along with their CLIP scores in Figure 6. High-value samples align with higher CLIP scores, while low-value ones are more concentrated in the lower range (half below 0.2). Furthermore, we calculate their correlation on random samples. We normalize both scores into the range 0 to 1 for clarity The observed positive correlation supports mimic scores in identifying high-quality samples.
Figure 5: The filtering logic behind Grad-Mimic aligns with handcrafted heuristics, particularly in caption length. Low-value samples (left figure) tend to have captions that are either too short or exceed the maximum token length, i.e., 77.
((a))
((b))
Figure 6: Mimic score distribution and correlation to CLIP score.
@@
We explore an interesting application of mimic score: predicting the pretraining dataset. Specifically, we ask: how accurately can Grad-Mimic check whether a given sample was used to train a reference model? We test this hypothesis on the small-scale DataComp dataset. We first apply various pre-built filters to curate datasets and train CLIP models on each. Then, we target their final-layer weights to mimic and apply Grad-Mimic to train on the entire data pool. Our goal is to see whether the top-ranked samples (matched in size to the curated datasets) identified by mimic score appear in their training dataset.
We evaluate our approach against random selection using Jaccard similarity and percentage of overlap. As shown in Table 5, mimic score-based selection identifies more samples used to train the reference model compared to random sampling. We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset by simply mimicking final-layer weights without direct access to their filtering steps.
@@ 6 Conclusion
We introduce Mimic Score, an efficient data quality metric that assesses samples’ utility in model training. We quantify samples by measuring the alignment between gradients and the vector toward pretrained reference weights. Using this, we develop Grad-Mimic to automate data selection. Empirically, Grad-Mimic effectively prioritizes samples for learning, improving data efficiency and model performance. Moreover, Mimic Score aids in creating effective filters and estimating dataset quality. Resulting filters outperform human-designed ones and exhibit a strong correlation with dataset quality and performance gains.
7 Impact Statements
Efficient data selection is increasingly important for training foundation models. We demonstrate Grad-Mimic’s effectiveness in identifying and removing low-value samples, which enhances data efficiency and resulting model performance. We do not foresee explicit negative social impacts arising from Grad-Mimic. However, if the reference model used to evaluate sample utility is suboptimal, there remains a risk of refining low-quality datasets (we have an analysis of the impact of reference model quality in Appendix C). We believe that careful selection and validation of this reference model, a key area of future research, will minimize this risk and enable the creation of improved datasets.
@@ 8 Acknowledgments
We thank Kwei-Herng Lai, Prasad Gabbur, Meng Cao, Ryan Jia, Russ Webb, Vedant Joshi, Albert Ge, John Cooper, Harit Vishwakarma, Dyah Adila, Changho Shin, Satya Sai Srinath Namburi, Zhiqi Gao, Avi Trost, and Gabe Orlanski for their helpful feedback and valuable discussion.
Appendix A Grad-Mimic Algorithm
We summarize the algorithm steps in Grad-Mimic for computing mimic scores and updating model weights. Our algorithm supports various training paradigms, including both supervised learning and self-supervised learning. We validate its effectiveness through controlled simulation experiments (Sec. 5.1) and large-scale training of CLIP models on web-crawled datasets (Sec. 5.2). Furthermore, our experiments encompass a range of training configurations, such as full-parameter fine-tuning, linear probing, and training models from scratch.
@@ Appendix A Grad-Mimic Algorithm
normalized mimic scores $\overline{m}_{i,t}$, a well-trained model $\theta_{T}$
@@
     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright compute the vector as a guide
@@
         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright compute per-sample gradient
         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|}$triangleright compute mimic score
@@
     $\overline{m}_{:,t}=\frac{e^{m_{:,t}/\tau}}{\sum_{j=1}^{b}e^{m_{j,t}/\tau}}$triangleright normalize mimic scores
@@
Appendix B Experimental Details
We provide more details about our training setups and computational resources.
B.1 Simulation Experiments
In the experiments of simulating mislabeled samples (Sec. 5.1), we fine-tune ViT-B/16 models pretrained on the ImageNet-21k dataset [51]. Each dataset is trained with a batch size of 32, a learning rate of 1e-4, the AdamW optimizer, and for 5 epochs. We evaluate Grad-Mimic using various temperature values $\tau$ (1.0, 0.9, 0.8, 0.7, 0.6, and 0.5) and report the performance at $\tau=0.5$ in the main paper. Results for other temperature values are presented in Table 7. To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets but with different random seeds for weight initialization. These experiments were performed on a Nvidia Tesla A100.
B.2 DataComp Experiments
We adopt the training setup from DataComp [49, 17] and train CLIP models from scratch using contrastive objective over image-caption pairs. Each CLIP model is trained for 5 epochs with a batch size of 4096. The total number of seen samples is 12.8M for the small-scale dataset and 128M for the medium-scale dataset. These experiments run on 8 Nvidia Tesla A100s.
We test Grad-Mimic using different parts of weights in the reference model, specifically the final MLP layer in the image and text encoders. Grad-Mimic’s performance is evaluated using 38 diverse downstream tasks [17] with various temperature values $\tau$ (0.03, 0.05, 0.07, 0.3, and 0.5). We report results at $\tau=0.5$ and $\tau=0.05$ in the main body in Table 3, but full results are presented in Table 6.
ImageNet
dist. shifts
VTAB Retrieval
Average over
38 datasets ($\uparrow$)
Vanilla Training 0.026 0.035 0.139 0.114 0.131
0.03 0.026 0.035 0.153 0.115 0.136
0.05 0.026 0.035 0.147 0.114 0.135
0.07 0.026 0.035 0.151 0.116 0.135
0.3 0.026 0.034 0.154 0.112 0.137
Last MLP Layer in Text Encoder 0.5 0.027 0.035 0.152 0.114 0.139
0.03 0.025 0.035 0.149 0.114 0.133
0.05 0.026 0.037 0.162 0.118 0.146
0.07 0.027 0.036 0.166 0.118 0.145
0.3 0.026 0.035 0.161 0.114 0.144
Vanilla Training 0.171 0.148 0.253 0.217 0.254
Medium Grad-Mimic
Last MLP Layer
in Image Encoder
0.05 0.169 0.151 0.262 0.216 0.258
Table 6: Full Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic consistently improves CLIP model performance across all temperature settings.
Appendix C Ablation Studies
@@
Table 7: Stage 1 Results in Simulation Experiment using Various Temperature: Lower temperature makes the model focus more on training with high-value samples, resulting in a higher testing accuracy.
@@
We test Grad-Mimic using different temperature values and compare them to baseline methods (Mini-batch SGD and Grad-Match [24]). The results in the simulation experiments are presented in Table 7. In these datasets, we set the noise level to 0.3 and fine-tune ViT-B/16 model under linear probing configuration. Grad-Mimic outperforms the baseline methods in all temperature settings. Lower temperatures yield better testing accuracy, as they normalize mimic scores in the way that encourages the model to focus more on high-value samples during training.
DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
Ref. Model 60.00 56.98 97.41 89.45 94.18 77.89 79.32
Mini-Batch SGD 51.91 47.71 44.44 28.64 21.50 15.43 96.26 95.49 93.56 87.33 85.55 83.51 92.86 92.14 91.19 75.56 74.38 73.25 72.09 69.46 66.40
Rho-Loss 54.10 52.39 48.46 41.55 35.71 30.02 97.10 96.83 96.49 88.53 87.79 86.86 94.07 93.83 93.82 76.82 76.05 75.93 75.36 73.77 71.93
Table 8: Grad-Mimic outperforms other reference model-based methods: Using the last-layer weights of reference model is more effective to prioritize samples for learning while reducing memory usage and computational overhead.
Comparison to Reference Model-Based Approaches.
We compare Grad-Mimic to Rho-Loss [15], a method that prioritizes samples for learning based on excess loss, computed as the difference between the loss on the current model $\ell(\theta_{t};s_{i})$ and the loss on the pretrained reference model $\ell(\theta_{\text{ref}};s_{i})$. We follow our simulation experiment setup and evaluate the performance of fine-tuned models on the testing datasets.
As shown in Table 8, Grad-Mimic consistently outperforms Rho-Loss across all datasets and noise levels. Furthermore, Grad-Mimic offers efficiency advantages. Unlike Rho-Loss, which requires loading the entire reference model and performing additional inference passes, Grad-Mimic computes mimic scores directly from the reference model’s weights. In our experiments, we demonstrate that even using only a subset of the weights (e.g., the last-layer weights) is sufficient, further reducing memory usage and improving computational efficiency.
Figure 7: Grad-Mimic can outperform influence function-based methods: Using model weights as a guide offer advantages on their effectiveness and greater accessibility.
Comparison to Influence Function-based Approaches.
Data influence functions provide a principled way to evaluating the impact of individual training samples on model parameters [52, 53], but their computational cost often limits practical use. Several studies have explored efficient approximations that estimate sample influence by measuring gradient alignment with a validation dataset [54, 55]. We adapt these methods within Grad-Mimic framework by replacing the vector induced by reference model’s weights ($v_{t}$) with gradients computed from validation dataset. We evaluate these adaptations on six image datasets by adding 50% label noise and assess the impact of validation set size by varying the number of samples drawn from each class (1, 5, 10, 15, and 20).
Results in Figure 7 show that while larger validation sets improve data efficiency by enabling model to focus on the influential samples, Grad-Mimic—using pretrained reference weights and their induced direction—consistently outperforms validation dataset-based influence approximations. Moreover, relying on a labeled validation set risk introducing an unwanted dependency. For example, in the out-of-domain Flowers102 dataset, achieving satisfactory performance required over 2000 correctly labeled validation samples—yet still fell short of Grad-Mimic’s results.
Figure 8: Grad-Mimic can be robust to the quality of reference model.
Impact of Reference Model Quality.
Finally, we investigate how the quality of the reference model impacts the effectiveness of Grad-Mimic, particularly in scenarios where only a weaker or distilled model is available. To simulate such scenarios, we inject zero-mean Gaussian noise $(0,\sigma^{2})$ into the reference model’s weights. By varying the variance of the added noise, we analyze the performance of models trained with Grad-Mimic against standard training methods that do not leverage a reference model.
We present this analysis in Figure 8. As expected, increasing noise variance degrades the reference model’s performance, which in turn lowers the effectiveness of Grad-Mimic. However, we observe that even under significant noise, models trained with Grad-Mimic still outperform Mini-batch SGD and can even surpass the performance of the noisy reference model itself. This highlights that while a high-quality reference model is ideal, even a weaker proxy can provide valuable guidance and improve training outcomes over traditional baselines.